The Metro Real Estate Association is preparing a pamphlet that they feel might be of interest to prospective home buyers in the Southlake and Northwood areas of the city. One item if interest is the length of time the seller occupied the home. A sample of 17 homes sold recently in Southlake revealed that the mean length of ownership was 9.2 years with a standard deviation of 2.9 years. A sample of 15 homes in Northwood revealed that the mean length of ownership was 7.4 years with a standard deviation of 2.1 years. Can we conclude that the Southlake residents owned their homes for a longer period of time compare to residents in the Northwood area? What is the decision rule? α = 0.05.

Group of answer choices

Reject Ho if test statistics > 1.697

Reject Ho if test statisitcs > 1.645

Reject Ho if test statistics > 2.042

Reject Ho if test statistics > 1.96
Answer #1

Sol:

Ho:mu1=mu2

Ha:mu1>mu2

alpha=0.05

df=n1+n2-2=17+15-2=30

p value in excel

=T.INV(0.05,30)

=1.697260887

Reject Ho,if t>1.697

answer:

Reject Ho if test statistics > 1.697